OpenStudy (anonymous):

how to find the sum from 0-inifinity (1/2^(k+3))

OpenStudy (anonymous):

\[\sum_{0}^{ \infty} 1/2^{k+3}\]

OpenStudy (anonymous):

find \[\sum_0^{\infty}\frac{1}{2^k}\] and subtract first 3 terms

OpenStudy (anonymous):

i got 1/4 could you verify that for me